OpenStudy (anonymous):

solve by completing the square x^2+5/2x=5

OpenStudy (anonymous):

is that (5/2)x or (5/2x)?

OpenStudy (anonymous):

I think you mean (5/2)x.....so you have x^2 + (5/2)x =5 To complete the square you need to divide the second term (5/2) by 2 = (5/4) and then square it which = (25/16). You then need to add that to both sides so you have x^2 +(5/2)x + (25/16) = 5 + (25/16)